OpenStudy (ksaimouli):

solve for x using quadratic formula x2+10x=9

OpenStudy (anonymous):

how about this? \[x^2+10x=9\] \[(x+5)^2=9+5^2\] \[(x+5)^2=34\] \[x+5=\pm\sqrt{34}\] \[x=-5\pm\sqrt{34}\]

OpenStudy (anonymous):

quadratic formula is harder to use here because it forces you to have a denominator. if the "middle term" has an even coefficient it is much easier to complete the square, which is really all the formula does for you anyway

OpenStudy (anonymous):

first you would have to write \[x^2+10x-9=0\] then use \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \[a=1,b=10,=-9\] you would get \[x=\frac{-10\pm\sqrt{136}}{2}\] and then you would have to simpify

OpenStudy (ksaimouli):

simpify please

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

when you simplify you get \[x=-5\pm\sqrt{34}\]

OpenStudy (ksaimouli):

ok

OpenStudy (anonymous):

oh you mean how how do you simplify \[x=\frac{-10\pm\sqrt{136}}{2}\]?

OpenStudy (ksaimouli):

ya

OpenStudy (anonymous):

i can show you if you like. you will get \[-5\pm\sqrt{34}\] but you have to be careful

OpenStudy (ksaimouli):

ok

OpenStudy (anonymous):

first of all you have to take care of \[\sqrt{136}\] is it \[\sqrt{4\times 34}=2\sqrt{34}\] because the square root of 4 is 2

OpenStudy (anonymous):

then you have \[\frac{-10\pm2\sqrt{34}}{2}\] factor out a 2 to get \[\frac{2(-5\pm\sqrt{34}}{2}\] then cancel to get \[-5\pm\sqrt{34}\]

OpenStudy (anonymous):

you can see why completing the square is easier here

OpenStudy (ksaimouli):

ya

OpenStudy (anonymous):

i don't have to do all that simplification at the end. get the answer in 4 easy steps

OpenStudy (ksaimouli):

thanks

OpenStudy (anonymous):

you can see the four steps at the top of this post. they are simpler than all this quadratic formula mes

OpenStudy (anonymous):

yw